OpenStudy (anonymous):

A line has a slope of 1/2 and a run = 50. What is the rise? 25 50 100

OpenStudy (eyust707):

Slope = rise/run

OpenStudy (anonymous):

slope = rise/run 1/2 = rise/50 how do you suppose you find the rise?

OpenStudy (anonymous):

so 50?

OpenStudy (anonymous):

50/50 =1 and not 1/2

OpenStudy (anonymous):

oh so 25/50=1/2